OpenStudy (anonymous):

Proving identities. Csc0/cot0 - cos0 = 1+ sin0/sin^30

OpenStudy (anonymous):

change cosec and cot is terms of sin and cos ..in LHS

OpenStudy (anonymous):

can u write the problem in the equation editor .?

OpenStudy (anonymous):

can you show me step by step how to solve this?

OpenStudy (anonymous):

please first ..present that question using equation editor.. the RHS part is not clear to me .

OpenStudy (anonymous):

its 1+sinu/sin^3u

OpenStudy (anonymous):

sin^3u is divided by 1+sinu or only sinu .??

OpenStudy (anonymous):

1+sinu is devided by sin^3u

OpenStudy (anonymous):

my equation editor is not working idk why sorry

OpenStudy (anonymous):

even mine is not working too .! change cosec as 1/sin ...and sot as cos/sin den put and simplify

OpenStudy (anonymous):

1/sinu is devided by cosu-cosu/sinu ???

OpenStudy (anonymous):

1/sinu divided by cosu/sinu - cosu

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

den (1/cosu) -cosu sin^2u/cosu

OpenStudy (anonymous):

is it devided by sin^2u/cosu??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is that all?

OpenStudy (anonymous):

no ..wondering what to do next ..to arrive at RHS

OpenStudy (anonymous):

hello i am awesome

OpenStudy (solomonzelman):

\[\frac{Cscθ }{Cotθ} - Cosθ = \frac{1+ Sinθ }{Sin^3θ}\] this is what you meant, here follow me,\[\huge\color{blue}{ \frac{Cscθ }{Cotθ} - Cosθ = \frac{1+ Sinθ }{Sin^3θ} } \] \[\huge\color{blue}{ \frac{ ~~~~\frac{1 }{Sinθ} ~~~~}{ ~~~\frac{Cosθ }{Sinθ}~~ } ~~ - Cosθ = \frac{1+ Sinθ }{Sin^3θ} } \] \[\huge\color{blue}{ \frac{1 }{Sinθ} \div \frac{Cosθ }{Sinθ} - Cosθ = \frac{1+ Sinθ }{Sin^3θ} } \] \[\huge\color{blue}{ \frac{1 }{Sinθ} \times \frac{Sinθ }{Cosθ} - Cosθ = \frac{1+ Sinθ }{Sin^3θ} } \] \[\huge\color{blue}{ \frac{1 }{Cosθ} - Cosθ = \frac{1+ Sinθ }{Sin^3θ} } \]

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \frac{1}{Cosθ} -Cosθ =\frac{1+Sinθ}{Sin^3θ}} \] \[\huge\color{blue}{ \frac{1}{Cosθ} - \frac{Cos^2θ}{Cosθ} =\frac{1+Sinθ}{Sin^3θ}} \] \[\huge\color{blue}{ \frac{1-Cos^2θ}{Cosθ} =\frac{1+Sinθ}{Sin^3θ}} \] \[\huge\color{blue}{ \frac{Sin^2θ}{Cosθ} =\frac{1+Sinθ}{Sin^3θ}} \]

OpenStudy (solomonzelman):

\[\huge\color{blue}{ Sin^5θ =(1+Sinθ)Cosθ} \]\[\huge\color{blue}{ Sin^5θ =Cosθ+SinθCosθ} \]

OpenStudy (solomonzelman):

too long